\begin{thebibliography}{32}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bird et~al.(2023)Bird, Ungless, and Kasirzadeh]{bird_typology_2023}
Charlotte Bird, Eddie Ungless, and Atoosa Kasirzadeh.
\newblock Typology of {Risks} of {Generative} {Text}-to-{Image} {Models}.
\newblock In \emph{Proceedings of the 2023 {AAAI}/{ACM} {Conference} on {AI}, {Ethics}, and {Society}}, pages 396--410, Montréal QC Canada, August 2023. ACM.
\newblock ISBN 9798400702310.
\newblock \doi{10.1145/3600211.3604722}.
\newblock URL \url{https://dl.acm.org/doi/10.1145/3600211.3604722}.

\bibitem[Bird(2020)]{BirdDecolonising2020}
Steven Bird.
\newblock Decolonising speech and language technology.
\newblock In \emph{Proceedings of the 28th International Conference on Computational Linguistics}, pages 3504--3519, Barcelona, Spain (Online), December 2020. International Committee on Computational Linguistics.
\newblock \doi{10.18653/v1/2020.coling-main.313}.
\newblock URL \url{http://dx.doi.org/10.18653/v1/2020.coling-main.313}.

\bibitem[Birhane et~al.(2022)Birhane, Kalluri, Card, Agnew, Dotan, and Bao]{birhane2022values}
Abeba Birhane, Pratyusha Kalluri, Dallas Card, William Agnew, Ravit Dotan, and Michelle Bao.
\newblock The {{Values Encoded}} in {{Machine Learning Research}}.
\newblock In \emph{2022 {{ACM Conference}} on {{Fairness}}, {{Accountability}}, and {{Transparency}}}, pages 173--184, Seoul Republic of Korea, June 2022. ACM.
\newblock ISBN 978-1-4503-9352-2.
\newblock \doi{10.1145/3531146.3533083}.

\bibitem[Bowker and Star(2000)]{bowkerSortingThingsOut2000}
Geoffrey~C. Bowker and Susan~Leigh Star.
\newblock \emph{Sorting Things out: Classification and Its Consequences}.
\newblock MIT Press, 2000.
\newblock ISBN 0-262-52295-0.

\bibitem[{boyd}(2024)]{boyd2024RisksVsHarms}
danah {boyd}.
\newblock Risks vs. {{Harms}}: {{Youth}} \& {{Social Media}}.
\newblock \emph{Data: Made Not Found (by danah)}, October 2024.

\bibitem[Chan et~al.(2023)Chan, Salganik, Markelius, Pang, Rajkumar, Krasheninnikov, Langosco, He, Duan, Carroll, Lin, Mayhew, Collins, Molamohammadi, Burden, Zhao, Rismani, Voudouris, Bhatt, Weller, Krueger, and Maharaj]{chan2023harms}
Alan Chan, Rebecca Salganik, Alva Markelius, Chris Pang, Nitarshan Rajkumar, Dmitrii Krasheninnikov, Lauro Langosco, Zhonghao He, Yawen Duan, Micah Carroll, Michelle Lin, Alex Mayhew, Katherine Collins, Maryam Molamohammadi, John Burden, Wanru Zhao, Shalaleh Rismani, Konstantinos Voudouris, Umang Bhatt, Adrian Weller, David Krueger, and Tegan Maharaj.
\newblock Harms from {Increasingly} {Agentic} {Algorithmic} {Systems}.
\newblock In \emph{2023 {ACM} {Conference} on {Fairness}, {Accountability}, and {Transparency}}, pages 651--666, Chicago IL USA, June 2023. ACM.
\newblock ISBN 9798400701924.
\newblock \doi{10.1145/3593013.3594033}.
\newblock URL \url{https://dl.acm.org/doi/10.1145/3593013.3594033}.

\bibitem[Cooper et~al.(2024)Cooper, Heldreth, and Hutchinson]{CooperIts2024}
Ned Cooper, Courtney Heldreth, and Ben Hutchinson.
\newblock ``{It’s how you do things that matters}'': Attending to process to better serve {Indigenous} communities with language technologies.
\newblock In \emph{Proceedings of the 18th Conference of the European Chapter of the Association for Computational Linguistics (Volume 2: Short Papers)}, pages 204--211, 2024.
\newblock URL \url{https://aclanthology.org/2024.eacl-short.19.pdf}.

\bibitem[Domínguez~Hernández et~al.(2024)Domínguez~Hernández, Krishna, Perini, Katell, Bennett, Borda, Hashem, Hadjiloizou, Mahomed, Jayadeva, Aitken, and Leslie]{dominguez_hernandez_mapping_2024}
Andrés Domínguez~Hernández, Shyam Krishna, Antonella~Maia Perini, Michael Katell, Sj~Bennett, Ann Borda, Youmna Hashem, Semeli Hadjiloizou, Sabeehah Mahomed, Smera Jayadeva, Mhairi Aitken, and David Leslie.
\newblock Mapping the individual, social and biospheric impacts of {Foundation} {Models}.
\newblock In \emph{The 2024 {ACM} {Conference} on {Fairness}, {Accountability}, and {Transparency}}, volume~7, pages 776--796, Rio de Janeiro Brazil, June 2024. ACM.
\newblock \doi{10.1145/3630106.3658939}.
\newblock URL \url{https://dl.acm.org/doi/10.1145/3630106.3658939}.

\bibitem[Douglas and Wildavsky(1983)]{douglas1983RiskCultureEssay}
Mary Douglas and Aaron Wildavsky.
\newblock \emph{Risk and {{Culture}}: {{An Essay}} on the {{Selection}} of {{Technological}} and {{Environmental Dangers}}}.
\newblock University of California Press, 1983.
\newblock ISBN 978-0-520-90739-3.
\newblock \doi{10.1525/9780520907393}.

\bibitem[Espinoza(2009)]{asveld_incommensurability_2009}
Nicolas Espinoza.
\newblock Incommensurability: {The} {Failure} to {Compare} {Risks}.
\newblock In \emph{The ethics of technological risk}. Earthscan London, 2009.
\newblock URL \url{https://api.taylorfrancis.com/content/books/mono/download?identifierName=doi&identifierValue=10.4324/9781849772990&type=googlepdf}.

\bibitem[Haraway(1988)]{haraway1988situated}
Donna Haraway.
\newblock Situated knowledges: {{The}} science question in feminism and the privilege of partial perspective.
\newblock \emph{Fem. Stud.}, 14\penalty0 (3):\penalty0 575--599, 1988.
\newblock ISSN 0046-3663.
\newblock \doi{10.2307/3178066}.

\bibitem[Hutchinson et~al.(2022)Hutchinson, Rostamzadeh, Greer, Heller, and Prabhakaran]{hutchinson2022evaluation}
Ben Hutchinson, Negar Rostamzadeh, Christina Greer, Katherine Heller, and Vinodkumar Prabhakaran.
\newblock Evaluation {{Gaps}} in {{Machine Learning Practice}}.
\newblock In \emph{2022 {{ACM Conference}} on {{Fairness}}, {{Accountability}}, and {{Transparency}}}, pages 1859--1876, Seoul Republic of Korea, June 2022. ACM.
\newblock ISBN 978-1-4503-9352-2.
\newblock \doi{10.1145/3531146.3533233}.

\bibitem[Hutiri et~al.(2024)Hutiri, Papakyriakopoulos, and Xiang]{hutiri_not_2024}
Wiebke Hutiri, Orestis Papakyriakopoulos, and Alice Xiang.
\newblock Not {My} {Voice}! {A} {Taxonomy} of {Ethical} and {Safety} {Harms} of {Speech} {Generators}.
\newblock In \emph{The 2024 {ACM} {Conference} on {Fairness}, {Accountability}, and {Transparency}}, pages 359--376, Rio de Janeiro Brazil, June 2024. ACM.
\newblock \doi{10.1145/3630106.3658911}.
\newblock URL \url{https://dl.acm.org/doi/10.1145/3630106.3658911}.

\bibitem[ImaniGooghari et~al.(2023)ImaniGooghari, Lin, Kargaran, Severini, Jalili~Sabet, Kassner, Ma, Schmid, Martins, Yvon, and Schütze]{ImaniGooghariGlot2023}
Ayyoob ImaniGooghari, Peiqin Lin, Amir~Hossein Kargaran, Silvia Severini, Masoud Jalili~Sabet, Nora Kassner, Chunlan Ma, Helmut Schmid, André Martins, François Yvon, and Hinrich Schütze.
\newblock {Glot500}: Scaling multilingual corpora and language models to 500 languages.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 1082--1117, Toronto, Canada, July 2023. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2023.acl-long.61}.
\newblock URL \url{https://aclanthology.org/2023.acl-long.61}.

\bibitem[Lamont(2012)]{lamont2012comparative}
Mich{\`e}le Lamont.
\newblock Toward a comparative sociology of valuation and evaluation.
\newblock \emph{Annual review of sociology}, 38:\penalty0 201--221, 2012.
\newblock \doi{10.1146/annurev-soc-070308-120022}.

\bibitem[Martin et~al.(2020)Martin, Prabhakaran, Kuhlberg, Smart, and Isaac]{martin2020ExtendingMachineLearning}
Donald Martin, Vinodkumar Prabhakaran, Jill Kuhlberg, Andrew Smart, and William~S. Isaac.
\newblock Extending the {Machine Learning Abstraction Boundary}: {A Complex Systems Approach} to {Incorporate Societal Context}.
\newblock \emph{arXiv preprint arXiv:2006.09663}, 2020.
\newblock \doi{10.48550/ARXIV.2006.09663}.

\bibitem[Mateescu and Elish(2019)]{mateescu2019ai}
Alexandra Mateescu and Madeleine Elish.
\newblock {AI} in context: the labor of integrating new technologies.
\newblock Technical report, Data \& Society Research Institute, New York, 2019.
\newblock URL \url{https://datasociety.net/wp-content/uploads/2019/01/DataandSociety_AIinContext.pdf.}

\bibitem[Moss et~al.(2021)Moss, Watkins, Singh, Elish, and Metcalf]{moss2021AssemblingAccountabilityAlgorithmic}
Emanuel Moss, Elizabeth~Anne Watkins, Ranjit Singh, Madeleine~Clare Elish, and Jacob Metcalf.
\newblock Assembling {{Accountability}}: {{Algorithmic Impact Assessment}} for the {{Public Interest}}, 2021.
\newblock URL \url{https://datasociety.net/wp-content/uploads/2021/06/Assembling-Accountability.pdf}.

\bibitem[OpenAI et~al.(2024)OpenAI, Achiam, Adler, Agarwal, Ahmad, Akkaya, Aleman, Almeida, Altenschmidt, Altman, Anadkat, Avila, Babuschkin, Balaji, Balcom, Baltescu, Bao, Bavarian, Belgum, and {...Barret Zoph}]{openai_gpt-4_2024}
OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia~Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, Red Avila, Igor Babuschkin, Suchir Balaji, Valerie Balcom, Paul Baltescu, Haiming Bao, Mohammad Bavarian, Jeff Belgum, and {...Barret Zoph}.
\newblock {GPT}-4 {Technical} {Report}, March 2024.
\newblock URL \url{http://arxiv.org/abs/2303.08774}.
\newblock arXiv:2303.08774 [cs].

\bibitem[Pratap et~al.(2023)Pratap, Tjandra, Shi, Tomasello, Babu, Kundu, Elkahky, Ni, Vyas, Fazel-Zarandi, Baevski, Adi, Zhang, Hsu, Conneau, and Auli]{PratapScaling2023}
Vineel Pratap, Andros Tjandra, Bowen Shi, Paden Tomasello, Arun Babu, Sayani Kundu, Ali Elkahky, Zhaoheng Ni, Apoorv Vyas, Maryam Fazel-Zarandi, Alexei Baevski, Yossi Adi, Xiaohui Zhang, Wei-Ning Hsu, Alexis Conneau, and Michael Auli.
\newblock Scaling speech technology to 1,{000+} languages.
\newblock \emph{arXiv [cs.CL]}, May 2023.
\newblock URL \url{http://arxiv.org/abs/2305.13516}.

\bibitem[Rakova and Dobbe(2023)]{RakovaAlgorithms2023}
Bogdana Rakova and Roel Dobbe.
\newblock Algorithms as social-ecological-technological systems: an environmental justice lens on algorithmic audits.
\newblock In \emph{Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency}, FAccT '23, page 491, New York, NY, USA, June 2023. Association for Computing Machinery.
\newblock ISBN 9798400701924.
\newblock \doi{10.1145/3593013.3594014}.
\newblock URL \url{https://doi.org/10.1145/3593013.3594014}.

\bibitem[Rubambiza et~al.(2024)Rubambiza, Sengers, Weatherspoon, and Liu]{rubambiza2024SeamWorkSimulacra}
Gloire Rubambiza, Phoebe Sengers, Hakim Weatherspoon, and Jen Liu.
\newblock Seam {{Work}} and {{Simulacra}} of {{Societal Impact}} in {{Networking Research}}: {{A Critical Technical Practice Approach}}.
\newblock In \emph{Proceedings of the {{CHI Conference}} on {{Human Factors}} in {{Computing Systems}}}, pages 1--19. ACM, 2024.
\newblock ISBN 9798400703300.
\newblock \doi{10.1145/3613904.3642337}.

\bibitem[Scott(2020)]{scott2020SeeingStateHow}
James~C. Scott.
\newblock \emph{Seeing like a State: How Certain Schemes to Improve the Human Condition Have Failed}.
\newblock Veritas {{Paperbacks Ser}}. Yale University Press, New Haven, 2020.
\newblock ISBN 978-0-300-25298-9.

\bibitem[Selbst et~al.(2019)Selbst, Boyd, Friedler, Venkatasubramanian, and Vertesi]{selbstFairnessAbstractionSociotechnical2019}
Andrew~D. Selbst, Danah Boyd, Sorelle~A. Friedler, Suresh Venkatasubramanian, and Janet Vertesi.
\newblock Fairness and abstraction in sociotechnical systems.
\newblock In \emph{Proceedings of the 2019 {{Conference}} on {{Fairness}}, {{Accountability}}, and {{Transparency}}}, pages 59--68, 2019.
\newblock ISBN 978-1-4503-6125-5.
\newblock \doi{10.1145/3287560.3287598}.

\bibitem[Shelby et~al.(2023)Shelby, Rismani, Henne, Moon, Rostamzadeh, Nicholas, {Yilla-Akbari}, Gallegos, Smart, Garcia, and Virk]{shelby2023sociotechnical}
Renee Shelby, Shalaleh Rismani, Kathryn Henne, {\relax Aj}ung Moon, Negar Rostamzadeh, Paul Nicholas, N'Mah {Yilla-Akbari}, Jess Gallegos, Andrew Smart, Emilio Garcia, and Gurleen Virk.
\newblock Sociotechnical {{Harms}} of {{Algorithmic Systems}}: {{Scoping}} a {{Taxonomy}} for {{Harm Reduction}}.
\newblock In \emph{Proceedings of the 2023 {{AAAI}}/{{ACM Conference}} on {{AI}}, {{Ethics}}, and {{Society}}}, pages 723--741, Montreal QC Canada, August 2023. ACM.
\newblock ISBN 9798400702310.
\newblock \doi{10.1145/3600211.3604673}.

\bibitem[Smart and Kasirzadeh(2024)]{smart_beyond_2024}
Andrew Smart and Atoosa Kasirzadeh.
\newblock Beyond model interpretability: socio-structural explanations in machine learning.
\newblock \emph{AI \& SOCIETY}, September 2024.
\newblock ISSN 0951-5666, 1435-5655.
\newblock \doi{10.1007/s00146-024-02056-1}.
\newblock URL \url{https://link.springer.com/10.1007/s00146-024-02056-1}.

\bibitem[Solaiman et~al.(2023)Solaiman, Talat, Agnew, Ahmad, Baker, Blodgett, Chen, Daum{\'e}, Dodge, Duan, Evans, Friedrich, Ghosh, Gohar, Hooker, Jernite, Kalluri, Lusoli, Leidinger, Lin, Lin, Luccioni, Mickel, Mitchell, Newman, Ovalle, Png, Singh, Strait, Struppek, and Subramonian]{solaimanEvaluatingSocialImpact2023a}
Irene Solaiman, Zeerak Talat, William Agnew, Lama Ahmad, Dylan Baker, Su~Lin Blodgett, Canyu Chen, Hal Daum{\'e}, Jesse Dodge, Isabella Duan, Ellie Evans, Felix Friedrich, Avijit Ghosh, Usman Gohar, Sara Hooker, Yacine Jernite, Ria Kalluri, Alberto Lusoli, Alina Leidinger, Michelle Lin, Xiuzhu Lin, Sasha Luccioni, Jennifer Mickel, Margaret Mitchell, Jessica Newman, Anaelia Ovalle, Marie-Therese Png, Shubham Singh, Andrew Strait, Lukas Struppek, and Arjun Subramonian.
\newblock Evaluating the {{Social Impact}} of {{Generative AI Systems}} in {{Systems}} and {{Society}}, 2023.

\bibitem[Suchman et~al.(1999)Suchman, Blomberg, Orr, and Trigg]{suchman1999reconstructing}
Lucy Suchman, Jeanette Blomberg, Julian~E. Orr, and Randall Trigg.
\newblock Reconstructing {{Technologies}} as {{Social Practice}}.
\newblock \emph{American Behavioral Scientist}, 43\penalty0 (3):\penalty0 392--408, November 1999.
\newblock ISSN 0002-7642, 1552-3381.
\newblock \doi{10.1177/00027649921955335}.

\bibitem[{Te Mana Raraunga}(2016)]{Te_Mana_Raraunga2016}
{Te Mana Raraunga}.
\newblock Our charter.
\newblock \url{https://www.temanararaunga.maori.nz/tutohinga}, April 2016.
\newblock URL \url{https://www.temanararaunga.maori.nz/tutohinga}.
\newblock Accessed: 2023-6-16.

\bibitem[Weidinger et~al.(2022)Weidinger, Uesato, Rauh, Griffin, Huang, Mellor, Glaese, Cheng, Balle, Kasirzadeh, Biles, Brown, Kenton, Hawkins, Stepleton, Birhane, Hendricks, Rimell, Isaac, Haas, Legassick, Irving, and Gabriel]{weidinger2022taxonomy}
Laura Weidinger, Jonathan Uesato, Maribeth Rauh, Conor Griffin, Po-Sen Huang, John Mellor, Amelia Glaese, Myra Cheng, Borja Balle, Atoosa Kasirzadeh, Courtney Biles, Sasha Brown, Zac Kenton, Will Hawkins, Tom Stepleton, Abeba Birhane, Lisa~Anne Hendricks, Laura Rimell, William Isaac, Julia Haas, Sean Legassick, Geoffrey Irving, and Iason Gabriel.
\newblock Taxonomy of {Risks} posed by {Language} {Models}.
\newblock In \emph{2022 {ACM} {Conference} on {Fairness}, {Accountability}, and {Transparency}}, pages 214--229, Seoul Republic of Korea, June 2022. ACM.
\newblock ISBN 978-1-4503-9352-2.
\newblock \doi{10.1145/3531146.3533088}.
\newblock URL \url{https://dl.acm.org/doi/10.1145/3531146.3533088}.

\bibitem[Winner(1986)]{winnerWhaleReactorSearch1986}
Langdon Winner.
\newblock \emph{The {{Whale}} and the {{Reactor}}: {{The Search}} for {{Limits}} in a {{Technological Age}}}.
\newblock University of Chicago Press, Chicago, 1986.

\bibitem[Wynne(1992)]{wynne1992UncertaintyEnvironmentalLearning}
Brian Wynne.
\newblock Uncertainty and environmental learning: {{Reconceiving}} science and policy in the preventive paradigm.
\newblock \emph{Global Environmental Change}, 2\penalty0 (2):\penalty0 111--127, July 1992.
\newblock ISSN 09593780.
\newblock \doi{10.1016/0959-3780(92)90017-2}.

\end{thebibliography}

\documentclass{article}

\PassOptionsToPackage{numbers, compress}{natbib}
\usepackage[preprint]{neurips2024}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         

\title{Troubling Taxonomies in GenAI Evaluation}

\author{%
  Glen Berman\\
  Australian National University\\
  Canberra, ACT, Australia\\
  \texttt{glen.berman@anu.edu.au}\\
  \And
  Ned Cooper\\
  Australian National University\\
  Canberra, ACT, Australia\\
  \texttt{edward.cooper@anu.edu.au}\\
  \And
  Wesley Hanwen Deng\\
  Carnegie Mellon University\\
  Pittsburgh, Pennsylvania, USA\\
  \texttt{hanwend@andrew.cmu.edu}\\
  \And
  Ben Hutchinson\\
  Google Research\\
  Pyrmont, NSW, Australia\\
  \texttt{benhutch@google.com}\\
}

\begin{document}

\maketitle


\begin{abstract}
    To evaluate the societal impacts of GenAI requires a model of how social harms emerge from interactions between GenAI, people, and societal structures. Yet a model is rarely explicitly defined in societal impact evaluations, or in the taxonomies of societal impacts that support them. In this provocation, we argue that societal impacts should be conceptualised as application- and context-specific, incommensurable, and shaped by questions of social power. Doing so leads us to conclude that societal impact evaluations using existing taxonomies are inherently limited, in terms of their potential to reveal how GenAI systems may interact with people when introduced into specific social contexts. We therefore propose a governance-first approach to managing societal harms attended by GenAI technologies.
\end{abstract}

\section{Models of the social world in GenAI evaluations}

Recent scholarship has started to make explicit the normative values and commitments of GenAI and Machine Learning (ML) practices \cite{birhane2022values} and evaluation \cite{hutchinson2022evaluation}. 
In this provocation, we extend this line of inquiry by arguing that we need to attend to the implicit values and assumptions reflected in how \textit{societal impacts} are conceptualised and constructed through ML evaluations.\footnote{Societal `impacts' is the phrasing generally adopted in responsible AI literature, which we follow. `Impact', however, is suggestive of immediacy and collisions. (Mateescu and Elish have made a similar point about `deploy' \cite{mateescu2019ai}.) Social `outcomes' may be preferable. This encourages thinking about long-term and second- and third-order outcomes of introducing AI systems into society.} Doing so reveals that the work of assessing and managing societal impacts of GenAI is best conceptualised through a governance, rather than a prediction, frame.\looseness=-1

Evaluating GenAI's societal impacts requires a model of how these impacts manifest \cite{martin2020ExtendingMachineLearning}. This model, often implicit, enables interpretation of how GenAI systems interact with people and social structures; the model constructs particular social factors as capable of, and deserving of, measurement. We ask: what is the model of societal impacts reflected in existing efforts to evaluate GenAI systems? One avenue to understand this model is to look at taxonomies of societal impacts \cite[\textit{e.g.},][]{solaimanEvaluatingSocialImpact2023a, shelby2023sociotechnical, weidinger2022taxonomy} which provide conceptual infrastructure for societal impact evaluations of GenAI \cite[for an alternate approach see][]{rubambiza2024SeamWorkSimulacra}. We note that some taxonomies refer to societal risks or harms, rather than impacts. Whilst we are wary of conflating these terms \cite{boyd2024RisksVsHarms}, in this provocation, we use ``taxonomies of societal impacts'' as a catchall category for efforts to develop a classification or categorisation system for relating GenAI to social phenomena.

\section{Taxonomies of societal impacts}

Taxonomies of societal impacts exist for a range of GenAI technologies and components, including foundation models \cite{dominguez_hernandez_mapping_2024}, text-to-image models \cite{bird_typology_2023}, large language models \cite{weidinger2022taxonomy}, speech generation models \cite{hutiri_not_2024}, AI agents \cite{chan2023harms}, and GenAI or algorithmic systems more generally \cite{solaimanEvaluatingSocialImpact2023a, shelby2023sociotechnical}. These tools direct the attention of GenAI evaluators, and provide structure for GenAI evaluations, in at least three ways. First, taxonomies of societal impacts enable researchers and practitioners to think systematically about the potential consequences of deploying GenAI technologies. Taxonomy development is, therefore, ontological work that has far-reaching consequences for the way researchers and practitioners understand the relationship between GenAI technologies, people, and society \cite{moss2021AssemblingAccountabilityAlgorithmic, bowkerSortingThingsOut2000}. 
Indeed, the development of bespoke taxonomies for different GenAI technologies implies a conceptualisation of societal impacts that centres technology as the primary causal determinate of harms and positions technology developers as the critical actors in impact evaluations (for contrast, imagine a range of taxonomies for different social contexts).\looseness=-1

Second, taxonomies of societal impacts direct attention by navigating the tension between abstraction and contextualisation, which is present throughout AI development \cite{selbstFairnessAbstractionSociotechnical2019}. GenAI components tend to be understood within an abstract space that disregards the social context of their development \cite{martin2020ExtendingMachineLearning}. Social harms, meanwhile, are understood as being deployment, \emph{i.e.}, context, dependent \cite[\textit{e.g.},][]{solaimanEvaluatingSocialImpact2023a}. \citet{shelby2023sociotechnical}, in their taxonomy, attempt to navigate this tension by distinguishing between harms that originate with computational components of AI systems, and harms that originate in their deployment.
Yet, GenAI systems (\textit{e.g.}, ChatGPT) are deployed extremely broadly, cutting across vast swathes of different social contexts. In this context, taxonomies direct attention towards forms of harm that can be tested or detected at the abstract level of the GenAI model, within the broader software evaluation paradigm in which GenAI deployment occurs \cite[\textit{e.g.}, GPT4,][]{openai_gpt-4_2024}.

Third, taxonomies of societal impacts also direct attention towards prediction and trade-offs, centring the site of GenAI development rather than application or deployment contexts. Taxonomies are often framed as predictive tools, enabling practitioners to forecast risks of GenAI deployment \cite[\textit{e.g.},][]{weidinger2022taxonomy, bird_typology_2023}. Yet, in attempting to develop an exhaustive schema of potential impacts, taxonomies invite comparison, and trade-offs, across disparate categories of societal impacts \cite{douglas1983RiskCultureEssay}, 
such as ``trust in media and information'', ``community erasure'', and ``intellectual property and ownership'' \cite{solaimanEvaluatingSocialImpact2023a}. 
Implicit in this organisation is a conceptualisation of societal impacts as modular, independent, and commensurable, with GenAI developers positioned as arbitrators in determining which impacts to address, and how.\looseness=-1

\section{Conceptualising the societal impacts of GenAI}

A conceptualisation of societal impacts that centres GenAI technologies and GenAI developers may be useful, in terms of producing a discourse on social outcomes that is tractable within GenAI, but should be approached with caution. Critical questions to consider include: what factors should be centred when thinking about GenAI's societal impacts? what are the limits of societal impact prediction? how should evaluators balance different forms of societal impacts? To begin responding to these questions, we offer three premises for rethinking the relationship between GenAI, people, and society.\looseness=-1

Societal impacts should be understood as application- and context-specific \cite{suchman1999reconstructing, martin2020ExtendingMachineLearning} and indeterminate \cite{wynne1992UncertaintyEnvironmentalLearning}. Failure to do so produces an understanding of societal impacts that is universalising and self-fulfilling; the work of evaluating societal impacts becomes the work of extending patterns of social relations from one place to many \cite{haraway1988situated}. Societal impacts of GenAI should be thought of at the system level, with the GenAI system situated in a particular social context \cite[cf., regarding model explanations,][]{smart_beyond_2024}. Impacts manifest when a model is integrated into a sociotechnical system, and implemented in a specific social setting \cite{shelby2023sociotechnical}. Context-specificity makes predictions about societal impacts difficult and unverifiable. Consequently, taxonomies of societal impacts are inherently partial, always incomplete.\looseness=-1

Some societal impacts should be understood as incommensurable \cite{asveld_incommensurability_2009}. The scaling of large multilingual models to include many languages, including Indigenous languages, illustrates this phenomenon \cite{PratapScaling2023,ImaniGooghariGlot2023}. Such scaling is motivated by the assumption that language technologies should be accessible to everyone in their first language \cite{BirdDecolonising2020}, which leads to model evaluations focused on identifying and rectifying performance disparities across languages. Yet, how should evaluators reconcile issues of disparate performance with issues of Indigenous data sovereignty, given one strategy to improve GenAI performance is to collect more language data? Navigating these trade-offs is particularly problematic when the objectives of GenAI developers may diverge from those of local communities. Some Australian Aboriginal and Māori communities
prioritise managing cultural knowledge, including language data, to support intergenerational transmission rather than expanding access to language technology \cite{CooperIts2024,Te_Mana_Raraunga2016}. In contexts like these, while taxonomies of societal impacts can help GenAI practitioners identify a broad range of impacts, GenAI practitioners are not well-positioned to balance competing impacts--these are value-laden decisions that require community leadership. Taxonomies can support such leadership by enabling practitioners to identify relevant stakeholders associated with different societal impacts \cite{bird_typology_2023}.

Finally, questions of societal impacts are questions about social power. Taxonomies of societal impacts enable evaluators to decide what to include (and exclude) in their evaluations. This legitimises particular concerns or forms of impact as salient to GenAI. As evaluation practices mature and become standardised, they gain efficacy, in terms of their capacity to enforce the values, simplifications, and assumptions they reflect \cite{lamont2012comparative, scott2020SeeingStateHow}. The dominance of cost-benefit analysis in environmental impact evaluation, for example, supports a capitalist and extractive epistemology, in which the worth of the environment is expressed in monetary terms \cite{winnerWhaleReactorSearch1986}. 
Efforts to standardise societal impact evaluation---worthy as they are---should, therefore, be understood as sociopolitical efforts that can reify or resist particular social orders. Who determines which societal impacts to focus on matters.

\section{For a governance-first approach}

The conceptualisation of societal impacts sketched above suggests redirecting efforts away from evaluations of potential harms and towards a governance-first approach to GenAI oversight. If societal harms are contextually contingent and indeterminate, then anticipatory evaluations may not be as effective at identifying and mitigating impacts as robust governance and monitoring of GenAI deployment led by stakeholders or governments.
Reflecting this, a governance-first approach would demand accountability to, participation of, and deliberation within, stakeholders or communities impacted by GenAI deployments \cite{RakovaAlgorithms2023}--for example, to determine how to balance disparate impacts. While practical recommendations for implementing a governance-first approach are beyond the scope of this provocation, a helpful starting place is to reconsider the role of taxonomies in RAI evaluations. Taxonomies and other evaluation tools can serve as useful inputs to robust governance and accountability processes \cite{moss2021AssemblingAccountabilityAlgorithmic}. However, without first establishing sustainable, representative governance structures (or engaging with those that already exist), these tools risk generalising predictions of harms across diverging contexts, equating incommensurable impacts, and ultimately serving the interests of GenAI researchers and developers rather than affected communities.\looseness=-1

\begin{ack}
We thank Jochen Trumpf, Kathy Reid, and Timothy Neale for timely literature suggestions, and the anonymous reviewers for their thoughtful feedback. Glen Berman and Ned Cooper are supported by Australian Government Research Training Program (RTP) Scholarships.
\end{ack}

\bibliographystyle{plainnat}

\bibliography{references}

\end{document}